OpenStudy (anonymous):

evaluate the given expression for x=-6 2x 2 - 2x -7

OpenStudy (anonymous):

\[\Large \begin{array}{l} 2{x^2} - 2x - 7\\ = 2{( - 6)^2} - 2( - 6) - 7\\ = 2 \cdot 36 + 12 - 7\\ = 77 \end{array}\]

OpenStudy (anonymous):

Answer is 77 as said above but i wonder why do u guys post this type of questions im middle of many interesting questions?????

OpenStudy (anonymous):

Everyone's at a different level of mathematics Atchyut, and everyone needs help on different things. For someone not familiar with algebra, such a question as the above may be quite intimidating. So you can't say that it's out of place amidst questions on calculus or higher-levels of geometry or anything else, since it's all mathematics.

OpenStudy (anonymous):

sorry, Im new, where should I post, a friend told me about this site?

OpenStudy (anonymous):

Its k fine I'm sorry I didnt mean it actually I thought someone is playing and sorry guys please do understand and its my first day to this site too and so k

OpenStudy (anonymous):

Posting here is fine Dee2012, as long as it's related to mathematics ;)

OpenStudy (anonymous):

Well welcome to OpenStudy! =)

OpenStudy (anonymous):

is there a "lower level" forum, and yes this is my first algebra class.

OpenStudy (anonymous):

OK- thanks

OpenStudy (mr.math):

@Atchyut: If you want to solve some difficult problems, you can join the group "melta-math". The group is not very active, but there are some tough problems you can find there. fewscrewsmissing is invited as well.